Mike and Dave play the following game. Mike picks from the deck of cards. If he draws a heart, Dave gives him $5. If not he gives Dave $2.
A. Determine mikes expectation
B. Determine Daves expectation

Answer by stanbon(75887)

Mike and Dave play the following game. Mike picks from the deck of cards. If he draws a heart, Dave gives him $5. If not he gives Dave $2.
----
Mikes random number gain: 5, -2
A. Determine mikes expectation
M[E(x)] = (1/4)5 + (3/4)-2 = 5/4-6/4 = -1/4 or -$0.25
---------------------------------------------------------
B. Determine Daves expectation
Daves random number gain: 2 , -5
D[E(x)] = (3/4)2 + (1/4)(-5) = 6/4 -5/4 = +$0.25
===================================================
Cheers,
Stan H.
===================================================